A random sample of size 16 is drawn from a normal distribution with σ=9.0 for the purpose of testing:

H0:μ=30 versus
H1:μ≠30.

The experimenter chooses to define the critical region C to be the set of samples means lying in the interval (23.9,30.1).

What level of significance does the test have?
Answer #1

Interval = (23.9, 30.1)

Therefore sample mean = (23.9+30.1)/2 = 27

Margin of error = 30.1-27 = 3.1

Now,